Exactly how many degrees are there between the hands of a clock at 3:40 pm
Answer by edjones(8007)

The hr hand is 2/3 the way to the 4 or 11/36 around the clock.
11/36 * 360 = 110 deg
The minute hand is 8/12 around the clock.
8/12 * 360 = 240 deg
.
240-110=130 degrees between the hands.
.
Ed